# Trackable Agent-based Evolution Models at Wafer Scale

### v2

@@ Conclusion
Computing hardware with transformative capabilities is coming to market right now. This fact presents an immediate opportunity to bring orders-of-magnitude greater simulation scale to bear on grand challenges in artificial life. It is not unreasonable to anticipate possibility that with such resources some aspects of these open questions will be revealed to harbor more-is-different dispositions, in which larger scales demonstrate qualitativey different dynamics (anderson1972more). Riding the coattails of AI-workload-driven hardware development, itself largely driven by profound more-is-different payoffs in deep learning, provides perhaps the most immediate means toward this possibility.
Such an endeavor is a community-level challenge that will require extensive collaborative effort and pursuit of outside resources. Work presented here is an early step in methods and wide-ranging infrastructure development necessary to scale up what’s possible in digital evolution research. We have demonstrated new algorithms and software for phylogeny-enabled agent-based evolution on a next-generation HPC accelerator hardware platform. Microbenchmarking results show that proposed instrumentation algorithms achieve several-fold improvement in computational efficiency. Related work shows new algorithms to improve reconstruction quality in some cases, too (moreno2024guide). Benchmarks confirm that, including tracking operations, simple models at wafer scale can achieve quintillions of replications per day.
Special characteristics set agent-based digital evolution apart from many other HPC application domains and position it as a potentially valuable testbed for innovation and leadership. Among these factors are challenging workload heterogeneity (varying within a population and over evolutionary time), resiliency of global state to local perturbations, and perhaps unparalleled freedom to recompose underlying simulation semantics to accommodate hardware capabilities. Indeed, artificial life and digital evolution have played host to notable boundary-pushing approaches to computing at the frontiers of computing modalities such as best-effort computing, reservoir computing, globe-spanning time-available computing, and modular tile-based computing in addition to more traditional cluster and GPU-oriented approaches (moreno2021conduit; ackley2020best; ackley2023robust; heinemann2008artificial; miikkulainen2024evolving). Work done to scale up digital evolution simulation should be done with an eye for contributions back to broader HPC constituencies.
@@
The immediate goal of our presented work, however, is to build means for progress toward existing research agendas held across the realms of digital evolution and artificial life. To this end, our work has prioritized production of reusable CSL and Python software tools ready to be harnessed by the broader research community. In particular, CSL code implementing the presented island-model GA is defined in a modularized and extensible manner to support drop-in customization to instantiate any fixed-length genome content and fitness criteria corresponding to research interests within the community. We look forward to collaboration in broader tandem efforts to harness the Cerebras platform, and other emerging hardware, in follow-on work.

### v3

@@ Conclusion
Computing hardware with transformative capabilities is presently coming to market. This fact presents an immediate opportunity to bring orders-of-magnitude greater simulation scale to bear on grand challenges in artificial life. It is not unreasonable to anticipate the possibility that with such resources some aspects of these open questions will be revealed to harbor more-is-different dispositions, in which larger scales reveal qualitatively different dynamics (anderson1972more). Riding the coattails of AI-workload-driven hardware development, itself largely driven by profound more-is-different payoffs in deep learning, provides perhaps the most immediate means toward this possibility.
Such an endeavor is a community-level challenge that will require significant resources and collaborative effort. Work presented here is an early step in methods and infrastructure development necessary to scale up what is possible in digital evolution research. We have demonstrated new algorithms and software for phylogeny-enabled agent-based evolution on a next-generation HPC accelerator hardware platform. Microbenchmarking results show that proposed instrumentation algorithms achieve several-fold improvement in computational efficiency. Related work shows new algorithms to improve reconstruction quality in some cases, too (moreno2024guide). Benchmarks confirm that, including tracking operations, simple models at wafer scale can achieve quintillions of replications per day. In future work, it will be necessary to move beyond proof-of-concept and explore the limits of these capabilities in the context of more demanding, interaction-intensive models.
Special characteristics set agent-based digital evolution apart from many other HPC application domains and position it as a potentially valuable testbed for innovation and leadership. Among these factors are challenging workload heterogeneity (varying within a population and over evolutionary time), resiliency of global state to local perturbations, and perhaps substantial freedom to recompose underlying simulation semantics to accommodate hardware capabilities. Indeed, artificial life and digital evolution have played host to notable boundary-pushing approaches to computing at the frontiers of computing modalities such as best-effort computing, reservoir computing, global-scale time-available computing, and modular tile-based computing in addition to more traditional cluster and GPU-oriented approaches (moreno2021conduit; ackley2020best; ackley2023robust; heinemann2008artificial; miikkulainen2024evolving; ray1995proposal). Work done to scale up digital evolution simulation should be done with an eye for contributions back to broader HPC constituencies.
@@
Our goal in this work is to build new capabilities that empower research agendas across the digital evolution and artificial life community. To this end, we have prioritized making our CSL and Python software easily reusable by other researchers. In particular, CSL code implementing the presented island-model GA is modularized and extensible for drop-in customization to instantiate any fixed-length genome content and fitness criteria. We look forward to collaboration in broader tandem efforts to harness the Cerebras platform, and other emerging hardware, in follow-on work.